Diffusion Morning Report 03/03/11

Ted Berg tberg at stanford.edu
Thu Mar 3 07:13:48 PST 2011


Good Morning All,
     Looking Ok this morning. Epi is the problem de jeur- handling. 
Thermco LTO pump out for repair. TEOS  has a leak. Ebeam compressor 
going out for repair. Tylan 4 wet ox a bit thin. Thermco 2 no TLC. No 
other major problems at this time.

					
	Date: 	3/3/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*yellow* 	*thin ox* 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*leak* 	*.25/30* 	** 	*2*
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*11.67/30* 	** 	**
Tylan 10-Nitride 	*yellow* 	*door close issue* 	*4.51/30* 	** 	**
Tylan 11-LTO 	*up* 	*a bit thin* 	*15.75/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*down* 	*bad pump-out for repair* 	** 	** 	**
AMAT Epi 	*down* 	*handling* 	** 	** 	*1*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor going out for repair* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*B on line 224 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110303/be105e87/attachment.html>


More information about the diffusion mailing list